OpenStudy (anonymous):

The graph of 2x + 3y = 12 is shifted down 3 units. What is the equation of the shifted graph? please explain

OpenStudy (anonymous):

2x+3y=9 the slope will be he same, but the y intercept will be 3 down, thus 12 - 3 =9

OpenStudy (anonymous):

The equation of a line is y=mx+b Where slope is m and b is the y intercept. In order to find what our equation needs to be we must first convert it into linear form (y=mx+b). We have 2x+3y=12 we want y=mx+b Using algebra: 2x+3y=12 3y=12-2x subtract 2x from both sides y=12/3 - (2x)/3 divide both sides by 3 y=4-2x/3 y=-2x/3+4 rewritten Which is what we want. m=-2/3 b=4 b is the location on the y axis so if we move the location of our intercept down by 3, we have 4-3=1=b Thus our final equation is y=-2x/3+1